At 20g of CaCO 3 were put into 10 litre flask and heated to 800 C, 30% of CaCO 3 remained unreacted at equilibrium. K p for decomposition of CaCO 3 will be?

At 20g of CaCO3 were put into 10 litre flask and heated to 800 C, 30% of CaCOremained unreacted at equilibrium. Kp for decomposition of CaCOwill be?

Grade:12

1 Answers

Bhavya
8 years ago
CaCO3 ----------> CaO + CO2(g)

0.2 mole 0 0

0.06 0.14 0.14

0.06= 0.2-x therefore x= 0.14

so Kp= p (CO2) only CO2is gas so its pressure is considered

Now,we know PV=nRT thus P=nRT/V

now p = 0.14*0.0821*1073/10
= 1.233